# CoAR: Concept Injection into Autoregressive Models for Personalized Text-to-Image Generation

### v1

@@ Introduction
Since CoAR leaves the pre-trained AR model parameters untouched, it fully leverages the strong multimodal understanding and generative capacity of autoregressive models, which naturally support zero-shot subject–style composition. This design enables training-free customization of arbitrary subjects in arbitrary styles simply by concatenating their respective context tokens, allowing for flexible, efficient, and scalable rendering without fine-tuning. To further ensure clean and controllable generation, we introduce an Identity Mask during inference, as shown in Fig.˜2(d), which explicitly restricts attention flow between subject and style tokens. This mechanism prevents semantic interference and ensures that each token type contributes distinctly and independently to the final generation.
Experiments
Experimental Details
@@
Figure 4: Re-contextualization with our CoAR on DreamBench (Ruiz et al. 2023) and StyleDrop (Sohn et al. 2023) datasets.
@@
Benchmark. We evaluate subject-driven customization using the DreamBench dataset (Ruiz et al. 2023), which contains 30 subjects (21 objects, 9 animals), each with 3–5 reference images and 25 test prompts. Objects are tested with 20 recontextualization and 5 property modification prompts, while animals use 10 recontextualization, 10 accessorization, and 5 property modification prompts. For stylization, we adopt the StyleDrop dataset (Sohn et al. 2023) and follow ZipLoRA (Shah et al. 2024), training with a single reference image per style. Additionally, we evaluate subject-style hybrid generation by combining DreamBench subjects with StyleDrop styles.
@@
Implementation Details. All experiments are run on a single H800 GPU using Lumina-mGPT-7B FP-SFT (Liu et al. 2024a) as the backbone. For subject-driven personalization, we insert one text and one image token per layer into the first 9 layers. For stylization with a single reference image, the same token configuration is applied to the first 3 layers. The model is trained with a learning rate of 1e-2 and batch size 1, keeping all other Lumina-mGPT hyperparameters unchanged. We set $\alpha$ = 1e-2, $\beta$ = 5e-4, and both $\lambda_{1}$ and $\lambda_{2}$ to 0.5. Each subject and style is trained for 1000 and 600 steps, respectively.
CoAR (Ours) B-LoRA ZipLoRA
@@
Table 2: Comparison results of our CoAR and two competitors on the style personalization task.
@@
Evaluation Metrics. For subject-driven customization tasks, we follow established protocols and employ CLIP-I (Radford et al. 2021) and DINO (Caron et al. 2021) for subject fidelity evaluation, along with CLIP-T (Radford et al. 2021) for prompt fidelity assessment. Specifically, CLIP-I computes the average cosine similarity between CLIP embeddings of real and generated images. DINO measures the average cosine similarity of ViT-S/16 DINO embeddings between real and generated images. CLIP-T evaluates the cosine similarity between CLIP text embeddings of prompts and CLIP image embeddings of generated outputs. For stylization tasks, we adopt metrics from ZipLoRA (Shah et al. 2024), which use CLIP features for both style and text alignment, and DINO features for subject alignment.
@@
Competitors. For the subject-driven personalization, we compare our approach with several baselines: Textual Inversion (TI) (Gal et al. 2022), DreamBooth (Ruiz et al. 2023), and two DreamBooth-based competitors (Wu et al. 2025), which share the same autoregressive backbone as our method. We also include two recent approaches specifically designed for multimodal autoregressive models (Sun et al. 2025; Wu et al. 2025). For the style personalization, we compare with recent diffusion-based methods: ZipLoRA (Shah et al. 2024) and B-LoRA (Frenkel et al. 2024).
Main Results
Quantitative Results of Subject-Driven Personalization.
@@
Table 1 shows the comparison results of our CoAR and the competitors on the DreamBench dataset. From the results, we have the following observations. First, CoAR achieves superior performance across multiple dimensions. In the CLIP-I metric, CoAR achieves a leading score of 0.8151, surpassing the second-best DreamBooth (Imagen) by 0.4%, demonstrating its exceptional image fidelity. For CLIP-T, CoAR sets a new benchmark at 0.3184, outperforming the second-best method, Proxy-Tuning by 0.29%, highlighting its superior cross-modal alignment capability. While CoAR ranks second on the DINO metric, it is substantially more efficient and user-friendly than the top-performing method, Proxy-Tuning, which requires training an additional diffusion model and generating hundreds to a thousand images per subject for data augmentation.
Quantitative Results of Style Personalization.
@@
Table 2 presents a comparison between our CoAR model and two baselines: B-LoRA (Frenkel et al. 2024) and ZipLoRA (SDXL) (Shah et al. 2024). The results demonstrate that CoAR outperforms B-LoRA across all three evaluation metrics, achieving the highest text-alignment score among all competitors. While CoAR trails ZipLoRA slightly in subject alignment (0.604 vs. 0.655), it achieves comparable style preservation (0.594 vs. 0.597), superior text alignment (0.308 vs. 0.272), and does so without requiring any additional training.
Qualitative Results of Subject-Driven Personalization.
@@
Fig.˜3 illustrates some generated samples of our CoAR compared with two most related competitors, i.e., PersonalAR (Sun et al. 2025) and Proxy-Tuning (Wu et al. 2025). From the results, we observe that images generated by CoAR exhibit superior prompt fidelity. For instance, given the instruction “wearing a black top hat and a monocle", neither PersonAR nor Proxy-Tuning successfully generates the “monocle", whereas our CoAR does. Moreover, our generated images better adhere to physical plausibility. In the second row of PersonAR, the generated vase and backpack are both floating, which is unrealistic, while such artifacts are absent in our results. More qualitative results are provided in the supplementary materials.
@@
Figure 5: Qualitative comparison of our CoAR and LoRA-based method for style personalization tasks.
Qualitative Results of Style Personalization.
@@
As shown in Fig.˜5, CoAR demonstrates superior performance compared to B-LoRA. While ZipLoRA preserves more subject details, such as color, CoAR exhibits stronger adherence to instructions when generating in novel scenes (e.g., “in the jungle” or “floating in water”). Additional qualitative results on style customization are presented in Fig.˜4, showing that CoAR effectively preserves the given style when generalizing the reference subject to novel scenes.
Ablation Studies
@@
Impacts of Different Losses. To assess the contribution of each loss component in our framework, we perform an ablation study by incrementally applying the LMCL, DPP, and CASR losses and evaluating performance on DINO, CLIP-I, and CLIP-T metrics, as shown in Tab. 3. Only applying LMCL yields baseline scores of 0.6610 (DINO), 0.7647 (CLIP-I), and 0.3096 (CLIP-T). Introducing DPP leads to consistent improvements across all metrics, increasing DINO to 0.6902 and CLIP-I to 0.7919. Although CLIP-T slightly decreases to 0.2968, the overall alignment benefits are clear. When CASR is further added, we observe the best overall performance: 0.7226 (DINO), 0.8151 (CLIP-I), and 0.3184 (CLIP-T), confirming that each loss component contributes positively to the alignment quality and subject fidelity.
@@
Impacts of Layer Numbers. To evaluate the impacts of depth of multimodal context token insertion, we insert tokens at layers 1, 3, 9, and 24. As illustrated in Fig.˜6. CLIP-T remains relatively stable across depths, while CLIP-I and DINO follow a rise-then-fall trend. Notably, inserting tokens at deeper layers (e.g., layer 24) leads to a decline in subject fidelity. We attribute this to overfitting caused by limited reference data, which becomes more pronounced as token influence shifts to higher-level semantic layers.
More experimental results are provided in the appendix.
@@
Conclusion
In this paper, we introduced CoAR, a novel and efficient framework for subject-driven and style-aware customization in multimodal autoregressive (AR) models. Unlike conventional fine-tuning or LoRA-based approaches, CoAR injects subject concepts directly into the pretrained distribution through a small number of learnable context tokens, avoiding any modification to pretrained parameters. Extensive experiments across personalization and stylization tasks demonstrate that CoAR achieves superior subject fidelity, instruction alignment, and style preservation while training fewer than 0.1M parameters, significantly more parameter-efficient than existing approaches. Moreover, our training-free strategy matches or surpasses modern LoRA-based diffusion methods, highlighting the potential of multimodal AR models for scalable and flexible personalization.
Appendix
Appendix A Training and Evaluation Details
Training Details. The training details of CoAR on different datasets are shown in Table˜4. For the style personalization task, only the provided single reference image is utilized. Consequently, the hyperparameters for Dual Prior Preservation (DPP) are not applied.
Evaluation Details. For the subject-style generation task, we selected 8 subjects from DreamBench (Ruiz et al. 2023), including 4 animals and 4 objects, as well as 6 styles from StyleDrop (Sohn et al. 2023). These yielded a total of 48 unique subject-style combinations, on which we evaluated our CoAR, ZipLoRA (Shah et al. 2024), and B-LoRA (Frenkel et al. 2024). For each combination, we assigned 10 distinct recontextualization prompts for both objects and living subjects, as illustrated in Table˜5.
@@ Appendix A Training and Evaluation Details
Table 4: Training settings of CoAR on DreamBench(Ruiz et al. 2023) and StyleDrop (Sohn et al. 2023) datasets.
subject context
objects in the snow
on the beach
with a city in the background
on top of a dirt road
on top of green grass
in the jungle
on the mountain
floating in water
on a picnic table
on top of a wooden floor
living subjects wearing a hat
with a crown
riding a bicycle
sleeping
in a boat
in the jungle
on the mountain
floating in water
on a picnic table
playing with a ball
@@
Appendix B Details of Parameter Calculation
For PersonalAR (Sun et al. 2025), we directly reference the parameter count reported in its original paper. As for Proxy-Tuning (Wu et al. 2025), since the paper does not provide explicit parameter statistics and the code is not publicly available, we provide a coarse estimate based on the hyperparameter configurations described in the paper.
Proxy-Tuning is a two-stage approach where a diffusion model is first trained, followed by an autoregressive (AR) model that is trained on data generated by the diffusion model. Both models are fine-tuned using Low-Rank Adaptation (LoRA) (Hu et al. 2022). In our estimation, the diffusion model is based on Stable Diffusion 3.5 (SD3.5 Large) (Esser et al. 2024), while the AR model is Lumina-mGPT 7B (Liu et al. 2024a). We estimate the total number of trainable LoRA parameters based on the typical LoRA application settings.
LoRA Parameter Calculation
@@ Appendix B Details of Parameter Calculation
$\text{LoRA params}=2dr$
@@
SD3.5 is based on the DiT (Peebles and Xie 2023) architecture, with 24 Transformer (Vaswani et al. 2017) blocks. Each block includes one self-attention and one cross-attention module. Assuming a hidden size $d=3072$, the LoRA (rank=64) parameters per attention module are:
@@
AR Model (Lumina-mGPT 7B)
@@
Total Parameter Count
@@
Appendix C Additional Experiments
Ablation
As shown in Fig.˜7, we first analyze the results using only Layer-wise Multi-modal Contextual Learning (LMCL), and then with the addition of the Dual Prior Preservation (DPP) loss. With the DPP constraint, the model better preserves the subject’s identity. However, when recontextualizing the subject in the scene “on a cobblestone street”, it still struggles to accurately render the background. In contrast, incorporating Context-Aware Self-Regularization (CASR) enables the model to successfully generalize to the new scene.
Fig.˜8 provides a visual illustration of the effect of our proposed CASR loss. When the loss weight is too small, the model tends to overfit: although subject fidelity remains high, generalization to novel contexts is limited. Conversely, when the weight is too large, the context tokens become under-optimized, leading to a significant degradation in subject fidelity.

### v2

@@ 3.6 Customize Any Subject in Any Style
Since DCoAR leaves the pre-trained AR model parameters untouched, it fully leverages the strong multimodal understanding and generative capacity of autoregressive models, which naturally support zero-shot subject–style composition. This design enables training-free customization of arbitrary subjects in arbitrary styles simply by concatenating their respective context tokens, allowing for flexible, efficient, and scalable rendering without fine-tuning. To further ensure clean and controllable generation, we introduce an Identity Mask during inference, as shown in Fig.˜2(d), which explicitly restricts attention flow between subject and style tokens. This mechanism prevents semantic interference and ensures that each token type contributes distinctly and independently to the final generation.
4 Experiments
4.1 Experimental Details
@@
Figure 4: Training-free subject-style generation with our DCoAR on DreamBench [ruiz2023dreambooth] and StyleDrop [sohn2023styledrop] datasets.
Figure 5: Re-contextualization with our DCoAR on DreamBench [ruiz2023dreambooth] and StyleDrop [sohn2023styledrop] datasets.
@@
Benchmark. We evaluate subject-driven customization using the DreamBench dataset [ruiz2023dreambooth], which contains 30 subjects (21 objects, 9 animals), each with 3–5 reference images and 25 test prompts. Objects are tested with 20 recontextualization and 5 property modification prompts, while animals use 10 recontextualization, 10 accessorization, and 5 property modification prompts. For stylization, we adopt the StyleDrop dataset [sohn2023styledrop] and follow ZipLoRA [shah2024ziplora], training with a single reference image per style. Additionally, we evaluate subject-style hybrid generation by combining DreamBench subjects with StyleDrop styles.
@@
Implementation Details. All experiments are run on a single H800 GPU using Lumina-mGPT-7B FP-SFT [lumina_mgpt] as the backbone. For subject-driven personalization, we insert one text and one image token per layer into the first 9 layers. For stylization with a single reference image, the same token configuration is applied to the first 3 layers. The model is trained with a learning rate of 1e-2 and batch size 1, keeping all other Lumina-mGPT hyperparameters unchanged. We set $\alpha$ = 1e-2, $\beta$ = 5e-4, and both $\lambda_{1}$ and $\lambda_{2}$ to 0.5. Each subject and style is trained for 1000 and 600 steps, respectively.
DCoAR (Ours) B-LoRA ZipLoRA
@@
Table 2: Comparison results of our DCoAR and two competitors on the style personalization task.
@@
Evaluation Metrics. For subject-driven customization tasks, we follow established protocols and employ CLIP-I [clip] and DINO [dino] for subject fidelity evaluation, along with CLIP-T [clip] for prompt fidelity assessment. Specifically, CLIP-I computes the average cosine similarity between CLIP embeddings of real and generated images. DINO measures the average cosine similarity of ViT-S/16 DINO embeddings between real and generated images. CLIP-T evaluates the cosine similarity between CLIP text embeddings of prompts and CLIP image embeddings of generated outputs. For stylization tasks, we adopt metrics from ZipLoRA [shah2024ziplora], which use CLIP features for both style and text alignment, and DINO features for subject alignment.
@@
Competitors. For subject-driven personalization on the DreamBench dataset, we compare our approach with several baselines. These include diffusion-based methods like Textual Inversion (TI) [textualinversion] and DreamBooth [ruiz2023dreambooth]. We also include a range of Unified Multimodal Models (MLLMs) that share an autoregressive architecture similar to ours. This category features DreamBooth adapted for MLLMs, Yo’Chameleon[yoChameleon], UniCTokens[UniCTokens], PersonalAR [personalAR], and Proxy-Tuning [wu2025proxytuning]. For style personalization, we compare with recent diffusion-based methods: ZipLoRA [shah2024ziplora] and B-LoRA [b-lora].
4.2 Main Results
@@
Quantitative Results of Subject-Driven Personalization. Table 1 shows the comparison results of our DCoAR and the competitors on the DreamBench dataset. From the results, we have the following observations. First, DCoAR achieves superior performance across multiple dimensions. In the CLIP-I metric, DCoAR achieves a leading score of 0.8151, surpassing the second-best DreamBooth (Imagen) by 0.4%, demonstrating its exceptional image fidelity. For CLIP-T, DCoAR sets a new benchmark at 0.3184, outperforming the second-best method, Proxy-Tuning by 0.29%, highlighting its superior cross-modal alignment capability. While DCoAR ranks second on the DINO metric, it is substantially more efficient and user-friendly than the top-performing method, Proxy-Tuning, which requires training an additional diffusion model and generating hundreds to a thousand images per subject for data augmentation.
@@
Quantitative Results of Style Personalization. Table 2 presents a comparison between our DCoAR model and two baselines: B-LoRA [b-lora] and ZipLoRA (SDXL) [shah2024ziplora]. The results demonstrate that DCoAR outperforms B-LoRA across all three evaluation metrics, achieving the highest text-alignment score among all competitors. While DCoAR trails ZipLoRA slightly in subject alignment (0.604 vs. 0.655), it achieves comparable style preservation (0.605 vs. 0.597), superior text alignment (0.308 vs. 0.272), and does so without requiring any additional training.
@@
Qualitative Results of Subject-Driven Personalization. Fig.˜3 illustrates some generated samples of our DCoAR compared with two most related competitors, i.e., PersonalAR [personalAR] and Proxy-Tuning [wu2025proxytuning]. From the results, we observe that images generated by DCoAR exhibit superior prompt fidelity. For instance, given the instruction “wearing a black top hat and a monocle", neither PersonAR nor Proxy-Tuning successfully generates the “monocle", whereas our DCoAR does. Moreover, our generated images better adhere to physical plausibility. In the second row of PersonAR, the generated vase and backpack are both floating, which is unrealistic, while such artifacts are absent in our results. More qualitative results are provided in the supplementary materials.
@@
Figure 6: Qualitative comparison of our DCoAR and LoRA-based method for style personalization tasks.
@@
Qualitative Results of Style Personalization. As shown in Fig.˜6, DCoAR demonstrates superior performance compared to B-LoRA. While ZipLoRA preserves more subject details, such as color, DCoAR exhibits stronger adherence to instructions when generating in novel scenes (e.g., “in the jungle” or “floating in water”). Additional qualitative results on style customization are presented in Fig.˜5, showing that DCoAR effectively preserves the given style when generalizing the reference subject to novel scenes.
4.3 Ablation Studies
@@
Impacts of Different Losses. We perform an ablation study by incrementally applying the LMCL, DPP, and CASR losses and evaluating performance on DINO, CLIP-I, and CLIP-T metrics, as shown in Tab. 3. Only applying LMCL yields baseline scores of 0.6610 (DINO), 0.7647 (CLIP-I), and 0.3096 (CLIP-T). Introducing DPP leads to consistent improvements across all metrics, increasing DINO to 0.7142 and CLIP-I to 0.8019. Although CLIP-T slightly decreases to 0.2977, the overall alignment benefits are clear. Alternatively, adding the CASR loss to the LMCL baseline improves all metrics, with scores of 0.7194 (DINO), 0.7905 (CLIP-I), and 0.3192 (CLIP-T). When all three losses (LMCL, DPP, and CASR) are applied together, we observe the best overall performance, achieving the highest scores for DINO (0.7226) and CLIP-I (0.8151), and a strong result for CLIP-T (0.3184). This confirms that each loss component contributes positively to the overall alignment quality and subject fidelity. As shown in Fig.˜8, we analyze the results using only Layer-wise Multi-modal Contextual Learning (LMCL), and then with the addition of the Dual Prior Preservation (DPP) loss. With the DPP constraint, the model better preserves the subject’s identity. However, when recontextualizing the subject in the scene "on a cobblestone street", it still struggles to accurately render the background. In contrast, incorporating Context-Aware Self-Regularization (CASR) enables the model to successfully generalize to the new scene.
@@
Impacts of Layer Numbers. To evaluate the impacts of depth of multimodal context token insertion, we insert tokens at layers 1, 3, 9, and 24. As illustrated in Fig.˜7. CLIP-T remains relatively stable across depths, while CLIP-I and DINO follow a rise-then-fall trend. Notably, inserting tokens at deeper layers (e.g., layer 24) leads to a decline in subject fidelity. We attribute this to overfitting caused by limited reference data, which becomes more pronounced as token influence shifts to higher-level semantic layers.
Impacts of Identity Mask. Figure˜9 demonstrates the critical role of our Identity Mask. Without it, the model suffers from severe concept contamination, causing the subject’s color to erroneously bleed into other scene elements. The mask prevents this by restricting attention flow between subject and style tokens, enforcing a clean disentanglement that ensures semantically coherent compositional generation. We provide more detailed analysis and results in the supplementary material.
Figure 9: Ablation results of Identity Mask.
5 Conclusion
In this paper, we introduced DCoAR, a framework that advances the concept-injection paradigm for AR model personalization by replacing previous shallow strategies with deep concept injection. Our method, Layer-wise Multimodal Context Learning, ensures robust feature propagation while keeping the backbone model completely frozen. Experiments demonstrate that DCoAR successfully bridges the fidelity gap, achieving the high performance while retaining superior parameter-efficiency. Moreover, its frozen backbone enables a training-free stylization strategy competitive with modern diffusion methods.
DCoAR: Deep Concept Injection into Unified Autoregressive Models for Personalized Text-to-Image Generation
Supplementary Material
6 Training and Evaluation Details
Training Details. The training details of DCoAR on different datasets are shown in Table˜4. For the style personalization task, only the provided single reference image is utilized. Consequently, the hyperparameters for Dual Prior Preservation (DPP) are not applied.
Evaluation Details. For the subject-style generation task, we selected 8 subjects from DreamBench [ruiz2023dreambooth], including 4 animals and 4 objects, as well as 6 styles from StyleDrop [sohn2023styledrop]. These yielded a total of 48 unique subject-style combinations, on which we evaluated our DCoAR, ZipLoRA [shah2024ziplora], and B-LoRA [b-lora]. For each combination, we assigned 10 distinct recontextualization prompts for both objects and living subjects, as illustrated in Table˜5.
@@ 6 Training and Evaluation Details
Table 4: Training settings of DCoAR on DreamBench[ruiz2023dreambooth] and StyleDrop [sohn2023styledrop] datasets.
Subject Context Subject Context
Objects in the snow Living Subjects wearing a hat
on the beach with a crown
with a city in the background riding a bicycle
on top of a dirt road sleeping
on top of green grass in a boat
in the jungle in the jungle
on the mountain on the mountain
floating in water floating in water
on a picnic table on a picnic table
on top of a wooden floor playing with a ball
@@
7 Details of Parameter Calculation
For PersonalAR [personalAR], we directly reference the parameter count reported in its original paper. As for Proxy-Tuning [wu2025proxytuning], since the paper does not provide explicit parameter statistics and the code is not publicly available, we provide a coarse estimate based on the hyperparameter configurations described in the paper. Proxy-Tuning is a two-stage approach where a diffusion model is first trained, followed by an autoregressive (AR) model that is trained on data generated by the diffusion model. Both models are fine-tuned using Low-Rank Adaptation (LoRA) [hu2022lora]. In our estimation, the diffusion model is based on Stable Diffusion 3.5 (SD3.5 Large) [sd3], while the AR model is Lumina-mGPT 7B [lumina_mgpt]. We estimate the total number of trainable LoRA parameters based on the typical LoRA application settings.
7.1 LoRA Parameter Calculation
@@
$\text{LoRA params}=2dr,$
@@
SD3.5 is based on the DiT [peebles2023scalablediffusionmodelstransformers] architecture, with 24 Transformer [transformer] blocks. Each block includes one self-attention and one cross-attention module. Assuming a hidden size $d=3072$, the LoRA (rank=64) parameters per attention module are:
@@
7.2 AR Model (Lumina-mGPT 7B)
@@
7.3 Total Parameter Count
@@
8 Additional Experiments
8.1 Ablation Study on Dual Prior Preservation (DPP)
To validate the design of our Dual Prior Preservation (DPP) strategy, we evaluate the individual contributions of the Class-based Next Token Prediction loss ($\mathcal{L}_{NTP_{cls}}$) and the KL Divergence constraint ($D_{KL}$). As shown in Table 6, applying these components in isolation leads to sub-optimal performance, highlighting the necessity of their joint application.
Specifically, employing only $\mathcal{L}_{NTP_{cls}}$ (Row 2) results in a significant drop in CLIP-T ($0.3192\rightarrow 0.3011$). This indicates that a hard reconstruction constraint without distributional regularization causes the model to overfit the visual appearance of the reference, leading to severe language drift and impaired re-contextualization capabilities.
Conversely, applying only $D_{KL}$ (Row 3) maintains text alignment but causes a sharp decline in subject fidelity (DINO drops to $0.7023$, CLIP-I to $0.7794$). This suggests that the soft distribution constraint alone is too conservative, preventing the model from learning the unique, fine-grained details of the specific subject.
However, when combined (Row 4), the two components exhibit a strong synergistic effect. The $\mathcal{L}_{NTP_{cls}}$ term ensures high subject fidelity (boosting CLIP-I to $\mathbf{0.8151}$), while the $D_{KL}$ term effectively regularizes the distribution to prevent language drift, maintaining a high CLIP-T score of $0.3184$. This confirms that the full DPP loss is essential for balancing identity preservation and text controllability.
Components Metrics
$\mathcal{L}_{NTP_{cls}}$ $D_{KL}$ DINO ($\uparrow$) CLIP-I ($\uparrow$) CLIP-T ($\uparrow$)
0.7194 0.7905 0.3192
0.7188 0.7893 0.3011
0.7023 0.7794 0.3189
0.7226 0.8151 0.3184
Table 6: Ablation study on the individual components of the Dual Prior Preservation (DPP) loss using the DreamBench dataset.
8.2 Visualization of the effects of CASR.
Figure˜10 provides a visual illustration of the effect of our proposed CASR loss. When the loss weight is too small, the model tends to overfit: although subject fidelity remains high, generalization to novel contexts is limited. Conversely, when the weight is too large, the context tokens become under-optimized, leading to a significant degradation in subject fidelity.